# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

### v2

@@
Abstract
Peer-to-peer (P2P) lending connects borrowers with lenders through online platforms but faces significant information asymmetry, as lenders often lack sufficient data to assess the creditworthiness of borrowers. This paper addresses this challenge by leveraging BERT (Bidirectional Encoder Representations from Transformers), a Large Language Model (LLM) known for its ability to understand contextual nuances in text, to analyze borrowers’ loan descriptions.
We apply transfer learning to make BERT distinguish between default and non-default loans using the loan descriptions of the Lending Club dataset. The resulting BERT-generated risk score demonstrates solid competence in discriminating between default and non-default loans. Furthermore, integrating the BERT-generated risk score with traditional variables enhances classifier performance, which demonstrates the complementary nature of advanced language model outputs in refining credit risk assessment methodologies.
However, the opacity of LLMs and potential biases highlight the need for transparent regulatory frameworks. This study opens new research avenues in P2P lending and AI, emphasizing the importance of trust and transparency in adopting advanced credit risk models.
keywords:
Credit Risk , P2P Lending , BERT , Transfer Learning , Explainable AI , XGBoost
[…]
[ucm] organization=Facultad de Informática, Universidad Complutense de Madrid, addressline=Calle del Prof. José García Santesmases, 9, city=Madrid, postcode=28040, state=Comunidad de Madrid, country=Spain
[…]
[itc] organization=Instituto de Tecnología del Conocimiento, Universidad Complutense de Madrid, addressline=Calle del Prof. José García Santesmases, 9, city=Madrid, postcode=28040, state=Comunidad de Madrid, country=Spain
{graphicalabstract}
{highlights}
Fine-tuned BERT successfully produces risk scores from loan descriptions.
Integrating BERT score with traditional variables boosts granting model performance.
The approach applies without manual annotation, reducing subjectivity and complexity.
Generating risk scores is instant with standard equipment.
Research is needed to enhance transparency and trust in these LLM-based approaches.
@@ 1 Introduction
Traditional credit scoring models do not harness the rich information embedded in the narratives submitted by loan applicants. Several approaches have tried to incorporate such information into credit risk models. The approaches include the extraction of linguistic metrics [3, 4], the use of topic modeling to characterize the underlying themes [5, 6], or a combination of both [7].
However, these methods have limitations in fully capturing the contextual nuances and semantic intricacies of loan descriptions. This study aims to address this gap by leveraging the capabilities of Large Language Models (LLMs), specifically BERT (Bidirectional Encoder Representations from Transformers) [8], which has shown exceptional performance in understanding and processing complex textual data.
BERT’s bidirectional training and ability to capture context at a granular level make it particularly suitable for tasks requiring deep semantic understanding. It has proven successful in fine-tuning for classification tasks [9], in particular domains such as biomedicine [10] and in specialized tasks like spam classification [11]. In this study, we apply transfer learning to train BERT to distinguish between default and non-default loans based on their textual descriptions.
Our work demonstrates BERT’s capability to effectively leverage descriptive data to generate a risk indicator that accurately classifies defaulted loans. Moreover, we show that integrating the BERT-generated risk score into traditional credit granting models can significantly enhance their predictive performance. This highlights BERT’s potential to improve the accuracy and reliability of credit risk models in P2P lending.
Furthermore, our findings highlight the need for greater transparency in adopting advanced models to build trust among users and regulatory entities. This study paves the way for future research, emphasizing the transformative potential of LLMs like BERT in the field of credit risk assessment.
@@ 2 Related work
More recently, Siering [7] investigated the effect of both aspects: topics and linguistic features. To extract topics, the author utilized a financial text analysis procedure [22] to build a domain dictionary. The identified topics describe the purpose of the loan, as well as the borrowers’ request for help, reliability, or appreciation. These topics were operationalized as binary indicator variables. The author employs text mining to create variables that measure aspects such as polarity, active orientation, readability, average sentence length, and word count. These variables are then used as input in a logistic regression. The results indicate that both linguistic and content-based factors contribute to predicting loan default probability, with the latter demonstrating greater significance. Analysis of variable contributions shows that certain factors increase or decrease default likelihood. Notably, expressions of reliability correlate positively with loan repayment.
Yet, a significant gap becomes apparent in applying state-of-the-art natural language processing techniques, including deep learning methods. In their work [6], Zhang et al. studied the transformer encoder’s ability to capture textual features from loan descriptions. These characteristics, alongside the traditional hard features derived from loan applications, were inputted into a neural network to predict the likelihood of loan default. The study highlights the effectiveness of transformers, showing better results when including textual loan descriptions compared to models that do not consider them. In this work, we explore using encoder-based LLMs, such as BERT, which has successfully improved classification in other fields.
3 An overview of Large Language Models and BERT
@@
To further elucidate the role of BERT in specialized applications, it is crucial to understand its capacity for transfer learning and fine-tuning. Transfer learning involves using a pre-trained model like BERT, which has initially learned general language patterns from a large corpus to a specific task or dataset. This technique allows us to take advantage of the rich linguistic representations without needing extensive computation from scratch. Fine-tuning involves adjusting the pre-trained model’s parameters to capture the nuances of the target task or application field by further training with new instances from the new context. For example, BioBERT is a BERT model fine-tuned for biomedical text mining tasks like named entity recognition and question answering [10]. Other adaptations have targeted text classification and sentiment analysis in specific datasets [9, 38].
Our research aims to develop a credit risk model by leveraging a BERT-based model. Specifically, we will fine-tune BERT for producing a predictive score indicative of the likelihood of loan default using the textual descriptions of loans provided by borrowers in their application forms. We will assess the scoring effectiveness in a granting model using the well-established dataset from the Lending Club P2P lending platform.
4 Description of the dataset
We have used a public data set of the P2P lending company Lending Club [^1] https://www.kaggle.com/wordsforthewise/lending-club, which is widely used in credit risk publications and the most widely used when dealing with the P2P market [39, 12]. However, instead of using the original dataset, which includes 2,260,699 loans granted by the company between 2007 and 2018, we will use a version modified for proposing granting models [40], used in [41, 42]. Since granting models determine which loans will be fully repaid, its estimation needs loans whose final status is known, that is, that were either fully repaid or defaulted. Thus, the dataset excludes loans in transitory states (in a grace period, late, etc.) and loans with no information on income and indebtedness which is essential to compute the input variables, resulting in 1,347,681 instances.
@@ 4 Description of the dataset
Table 4: Exploratory data analysis. Textual variable.
@@
5 Description of the methodology
This study assesses the enhancement of default prediction by integrating textual descriptions of loans preprocessed through an LLM. Specifically, we employ transfer learning and fine-tuning to enable the LLM to generate a score reflecting the probability of loan default, leveraging information derived from the loan’s textual description.
As our baseline, we employ a machine learning classifier that utilizes the complete set of variables available during the loan application process, encompassing both quantitative and categorical variables. Subsequently, we enhance the classifier by incorporating the LLM score—representing default risk inferred from processing the textual description with the LLM—as an additional input variable and proceed to evaluate the resulting differences. The data and code of our experiments are publicly available [^4] Link of the code will be included in final version..
The key components of the experiment are elaborated upon in this section.
@@ 5.1 Tuning the classifier
The classification algorithm used was XGBoost [43], which is trained using a stratified k-fold cross-validation with k being equal to 5 and where the instances are shuffled to avoid potential ordering biases in the dataset. A genetic algorithm [44] was employed to fine-tune the hyperparameters. The fitness value of each individual was the average balanced accuracy (BACC) in the 5 validation sets of the cross-validation. We use BACC as it accounts for the imbalanced nature of the dataset. In preliminary experiments, we also considered the area under the receiver operating characteristic (AUROC), which measures the model’s ability to discriminate between positive and negative examples regardless of the classification threshold chosen. However, we observed that the resulting XGBoost classifiers produced poor BACC values (similar to those from a naïve classifier that predicts the majority class) when using the standard 0.5 threshold to make the prediction. We also observed that XGBoost classifiers with extremely similar AUROC values produced very different results in terms of BACC. Thus, we decided to use the BACC measure as it resulted in classifiers with a more stable behavior.
In genetic optimization, each individual is characterized by its genes, that is, the considered hyperparameters of the XGBoost. Table 5 shows the hyperparameters together with their respective ranges.
@@
alpha 0.5 10
lambda 0.5 10
subsample 0.7 1
colsample_bytree 0.3 1
scale_pos_weight 0.1 10
n_estimators 2 500
The evolutionary strategy chosen for the optimization is the “Mu plus lambda” approach, denoted as $\mu+\lambda$, where $\mu$ represents the number of individuals to select for the next generation, and $\lambda$ indicates the number of children to produce at each generation. Unlike traditional approaches where children often replace parents, the $\mu+\lambda$ strategy involves adding both children and parents to produce the next generation. In the context of this research, we set $\mu$ to 150 and $\lambda$ to 150. This configuration, considering the addition of both parents and offspring, results in generations comprising a total of 300 individuals.
Initially, pairs of parents are chosen through tournament selection with a tournament size of 2. Subsequently, the children are generated employing a two-point crossover technique on the parents’ chromosomes with an 80% probability and applying a random resetting mutation with a 20% probability. The random resetting mutation implies that each gene of every child has a 20% chance of acquiring a new random value within its defined range. To create an offspring of 150 children, this selection, crossover, and mutation process is repeated 75 times. Finally, we combine the 150 children and the 150 parents, resulting in generations of 300 individuals, and select the top 150 ($\mu$) according to their fitness value to pass to the next generation.
The evolutionary process consists of 20 iterations, thereby generating a total of 3,000 individuals—each representing a distinct hyperparameter configuration. From this pool of configurations, the one exhibiting the highest fitness is ultimately chosen as the optimal outcome.
@@ 5.2 Generating a default score with BERT
In this section, we delineate the methodology employed to generate a default score based on the textual description of the loan. We initiate the process by applying transfer learning utilizing an LLM, specifically BERT in our case. The fine-tuned BERT model produces an outcome within the range of 0 to 1, offering a nuanced indicator rather than a binary classification. This subtle indicator is subsequently integrated into a classifier along with other input variables to predict the likelihood of loan default. The subsequent steps in this process are detailed next.
@@ 5.2.1 Transfer learning to produce the default score
As outlined in Section 3, during the transfer learning phase, we aim to exploit BERT’s advanced language understanding while minimizing the need to learn from scratch. To achieve this, we freeze the weights of all but the last hidden layer of BERT. This approach preserves the model’s pre-trained capabilities and prevents overfitting on our specific dataset, a phenomenon known as catastrophic forgetting [^6] A degradation issue during the retraining of pre-trained models, leading to loss of the original capabilities [45]..
Subsequently, in the fine-tuning stage, only the last BERT layer and the newly added layers are adjusted to better serve our specific task of generating a default score. These layers are optimized to enhance the model’s adaptability to our particular requirements, allowing slight parameter adjustments for improved task-specific performance, while maintaining the general language understanding gained from BERT’s initial pre-training. This strategy effectively balances specialized learning with the retention of valuable pre-trained knowledge.
We also explored various configurations for the extra layers to determine the one that produces the best results. Specifically, we explored the 126 configurations that result from combining the following options:
  • 1.
@@
The neural network is trained to predict the loan outcome (default or non-default) based on the textual description as input. As a loss function, we use the weighted binary cross-entropy, which measures the difference between the predicted probabilities and the true binary labels of the input data taking into account class weights [^9] Considering class weights is crucial given the imbalanced nature of our dataset.. The training was set up with a batch size of 64 and trained for 25 epochs, with an early stopper of 3 epochs.

### v3

@@
Abstract Peer-to-peer (P2P) lending connects borrowers and lenders through online platforms but suffers from significant information asymmetry, as lenders often lack sufficient data to assess borrowers’ creditworthiness. This paper addresses this challenge by leveraging BERT, a Large Language Model (LLM) known for its ability to capture contextual nuances in text, to generate a risk score based on borrowers’ loan descriptions using a dataset from the Lending Club platform. We fine-tune BERT to distinguish between defaulted and non-defaulted loans using the loan descriptions provided by the borrowers. The resulting BERT-generated risk score is then integrated as an additional feature into an XGBoost classifier used at the loan granting stage, where decision-makers have limited information available to guide their decisions. This integration enhances predictive performance, with improvements in balanced accuracy and AUC, highlighting the value of textual features in complementing traditional inputs. Moreover, we find that the incorporation of the BERT score alters how classification models utilize traditional input variables, with these changes varying by loan purpose. These findings suggest that BERT discerns meaningful patterns in loan descriptions, encompassing borrower-specific features, specific purposes, and linguistic characteristics. However, the inherent opacity of LLMs and their potential biases underscore the need for transparent frameworks to ensure regulatory compliance and foster trust. Overall, this study demonstrates how LLM-derived insights interact with traditional features in credit risk modeling, opening new avenues to enhance the explainability and fairness of these models.
Keywords: Credit Risk, Peer-to-Peer Lending, Natural Language Processing, BERT, Transfer Learning, Explainable AI
@@ 1 Introduction
Traditional credit scoring models often overlook the valuable information contained in loan applicants’ narratives [35]. Several methods have attempted to incorporate this data, including extracting linguistic metrics [16, 15], using topic modeling to identify underlying themes [55, 62], or combining both approaches [46]. These methods offer clear advantages, such as computational efficiency—particularly in the case of linguistic metrics—and greater interpretability. However, they also come with drawbacks, including limited ability to capture meaning—especially for linguistic metrics—, dependence on extensive preprocessing in the case of topic modeling, and an overall inability to fully grasp nuance and context.
In contrast, transformer-based models, such as BERT (Bidirectional Encoder Representations from Transformers) [13], offer a powerful alternative. While these models are less interpretable, they excel at understanding text by capturing semantic and contextual relationships at the word, sentence, and document levels. They also adapt flexibly to variations in style and structure. Moreover, leveraging pre-trained models is straightforward—by transferring general language knowledge from large corpora, these models can be fine-tuned for specific tasks or domains, achieving significant performance improvements with relatively little labeled data.
In sum, BERT’s bidirectional training and context-sensitive representations make it well-suited for tasks requiring deep semantic understanding. It has been successfully fine-tuned for various classification tasks [48], including applications in biomedicine [25] and specialized areas such as spam detection [52]. Recently, Xia et al. [57] demonstrated the effectiveness of a fine-tuned BERT model in discriminating P2P loans within the Chinese market.
In this study, we extend this line of research by applying BERT at the loan granting stage—a critical point where customer narratives hold greater importance due to the limited information available for decision-making. As our baseline method, we employ XGBoost, an efficient gradient-boosting algorithm that builds ensembles of decision trees to enhance performance and that has demonstrated its ability in loan granting scenarios [2]. We show that incorporating BERT-generated risk scores into a loan granting model significantly enhances predictive performance, and we delve deep into how BERT processes textual data and influences model behavior. Specifically, our analysis reveals that BERT captures a wide range of information from loan descriptions, including borrower-specific attributes, loan purposes, and linguistic features embedded in the text. Furthermore, we demonstrate that the inclusion of BERT-generated scores reshapes how credit models leverage other input variables, with the impact varying substantially across different loan purposes. While our findings highlight BERT’s potential to improve credit risk assessment in P2P lending, they also emphasize the importance of transparency in understanding what these models learn from text. Such transparency is crucial for building trust among stakeholders and ensuring acceptance by entities.
@@ 2 Related Work
2.1 Data Sources in Credit Risk Modeling: The Use of Loan Descriptions
@@ 2.2 Topic Modeling Approaches
Siering [46] recently examined the impact of both topical and linguistic features on loan default prediction. To extract topics, the author employed a financial text analysis method [29] to construct a domain-specific dictionary. The identified topics captured elements such as the loan purpose, the borrower’s requests for assistance, expressions of reliability, and appreciation. These topics were represented as binary indicator variables. Additionally, text mining techniques were used to generate features measuring attributes like polarity, active orientation, readability, average sentence length, and word count. These features were then incorporated into a logistic regression model, revealing that both linguistic and content-based factors contribute to predicting loan default probability, with content-based factors showing greater significance. The analysis further indicated that certain variables, such as expressions of reliability, positively correlate with loan repayment likelihood.
2.3 Advancements with Large Language Models
In recent years, several studies have begun to explore state-of-the-art natural language processing techniques, including deep learning models, for loan default prediction. In their work [62], Zhang et al. investigated transformer encoders for extracting textual features from loan descriptions. These features, combined with traditional hard features from loan applications, were input into a neural network to predict default probabilities. Their findings demonstrated the effectiveness of transformers, with models incorporating textual loan descriptions outperforming those that did not.
More recently, Xia et al. [57] explored the use of BERT-based LLMs to extract information from loan narratives, integrating these insights into both logistic regression and machine learning models like random forest and deep forest. While this work achieved improved predictive performance, it lacked transparency regarding the specific information captured by the LLM and how it influenced the classification model, leaving the results largely opaque and black-box.
2.4 Research Gap
Despite the progress made in leveraging textual data for credit risk modeling, several gaps remain. First, while advanced NLP techniques like BERT have been applied to loan descriptions, the interpretability of the extracted features and their influence on downstream models remains underexplored. For instance, a previous work [57] achieved improved performance but provided limited insights into how the BERT-derived features contributed to the model’s predictions. Second, there is a need for a more systematic integration of textual features with traditional financial variables in a way that enhances both predictive accuracy and interpretability.
Our study addresses these gaps by applying BERT to create a risk score from loan descriptions in the loan-granting process, a context where customer narratives play a crucial role. Unlike previous work, we analyze the score, including its relationship with other variables, its impact on the classification model, and on the results obtained across the predefined loan purposes segmented, aiming to provide more interpretability and insights into the decision-making process.
3 Basics of LLM architectures and BERT
@@
To further elucidate the role of BERT in specialized applications, it is crucial to understand its capacity for transfer learning and fine-tuning. Transfer learning involves using a pre-trained model like BERT, which has initially learned general language patterns from a large corpus to a specific task or dataset. This technique allows us to take advantage of the rich linguistic representations without needing extensive computation from scratch. Fine-tuning involves adjusting the pre-trained model’s parameters to capture the nuances of the target task or application field by further training with new instances from the new context. For example, BioBERT is a BERT model fine-tuned for biomedical text mining tasks like named entity recognition and question answering [25]. Other adaptations have targeted text classification and sentiment analysis in specific datasets [48, 17]. Finally, as already mentioned, [57] shows how fine-tuned Chinese BERT models can enhance the classification performance of default loans in the P2P market.
4 Dataset
We use a public data set of the P2P lending company Lending Club [^1] https://www.kaggle.com/wordsforthewise/lending-club, which is widely used in credit risk publications and the most widely used when dealing with the P2P market [1, 4]. However, instead of using the original dataset, which includes 2,260,699 loans granted by the company between 2007 and 2018, we use a version modified for proposing granting models [5], used in [2, 3]. Since granting models determine which loans will be fully repaid, its estimation needs loans whose final status is known (i.e., that were either fully repaid or defaulted). Thus, the dataset excludes loans in transitory states (in a grace period, late, etc.) and loans with no information on income and indebtedness, which is essential to compute the input variables, resulting in 1,347,681 instances.
@@ 4 Dataset
Table 4: Exploratory data analysis. Textual variable (desc).
@@
5 Methodology
This study employs transfer learning to enable an LLM to generate a score that reflects the likelihood of loan default based on textual descriptions. We explore how a fine-tuned LLM captures various aspects of loan descriptions that are indicative of default risk. Furthermore, we demonstrate that incorporating the LLM-generated score enhances the predictive accuracy of a loan-granting model and fundamentally alters how the model operates compared to when the score is absent.
Our baseline model is a machine learning classifier that utilizes all available variables from the loan application process, including both quantitative and categorical data. Specifically, we use XGBoost, which has been shown to deliver superior performance in similar loan-granting contexts [2]. To augment this baseline, we introduce the LLM-generated default risk score as an additional input feature. We analyze the information captured by this score and assess its impact on both prediction performance and the behavior of the resulting model.
The experimental setup is shown in Figure 1. Loan descriptions are processed by a fine-tuned BERT model, which outputs a BERT_score representing the probability of default. This score is then integrated with other input variables in the XGBoost classifier to produce the final prediction.
Figure 1: Diagram of the experiment architecture.
The key components of the methodology are elaborated upon in this section.
@@ 5.1 Tuning the classifier
The classification algorithm used was XGBoost [11], which is trained using a stratified $k$-fold cross-validation, dividing the dataset into $k$ subsets (here, $k=5$) and preserving the original class distribution in each fold to provide a more reliable evaluation in an imbalanced dataset as ours. The dataset is shuffled to eliminate biases derived from its original ordering and to ensure representative subsets, avoiding artificial patterns in the training. Furthermore, k=5 is used to balance the reduction of variance in the estimates with the need for sufficiently large partitions, a crucial factor in datasets with few instances and high heterogeneity as ours.
Furthermore, the instances are shuffled to avoid potential ordering biases in the dataset. To fine-tune the hyperparameters we used a genetic algorithm [20] to evolve candidate hyperparameter combinations and choose the one that maximizes the fitness measure, which was the average balanced accuracy (BACC) in the 5 validation sets of the cross-validation. We use BACC as it accounts for the imbalanced nature of the dataset. In preliminary experiments, we also considered the area under the receiver operating characteristic (AUROC), which measures the model’s ability to discriminate between positive and negative examples regardless of the classification threshold chosen. However, we observed that the resulting XGBoost classifiers produced poor BACC values (similar to those from a naïve classifier that predicts the majority class) when using the standard 0.5 threshold to make the prediction. We also observed that XGBoost classifiers with extremely similar AUROC values produced very different results in terms of BACC. Thus, we decided to use the BACC measure as it resulted in classifiers with more stable behavior.
In genetic optimization, each individual is characterized by its genes, that is, the considered hyperparameters of the XGBoost. We have included several kinds of parameters, including:
  • Parameters that adjust the sample weights, such as scale_pos_weight, which balance the weights of the classes and are useful in unbalanced datasets as ours.
  • Parameters that set up the behavior of the boosting algorithm, such as the learning rate (eta), the percentage of subsamples in each iteration (subsample), the number of learners (n_estimators), and the percentage of dataset features that use each learner (colsample_bytree).
  • Parameters that control the learning process of each tree, including its maximum depth (max_depth), regularization parameters (lambda and alpha), the loss reduction required to make a further partition on a leaf node (gamma), and the minimum number of weighted instances needed in a child node (min_child_weight).
Table 5 shows the hyperparameters together with their respective ranges, which were determined based on a combination of domain knowledge, preliminary experiments, and established practices in the literature [2, 61]. For instance, the learning rate (eta) was set between 0.001 and 0.5 to balance the trade-off between convergence speed and model performance. The max_depth parameter, which controls the maximum depth of a tree, was set between 2 and 12 to prevent overfitting while allowing sufficient model complexity. Similarly, the ranges for other hyperparameters, such as regularization terms and sampling rates, were selected to ensure a wide exploration of their effects, which are crucial for optimizing model performance in imbalanced datasets.
@@
The evolutionary strategy chosen for the optimization is the “Mu plus lambda” ($\mu+\lambda$) approach, where $\mu$ represents the number of individuals to select for the next generation, and $\lambda$ indicates the number of children to produce at each generation. Unlike traditional approaches where children often replace parents, the $\mu+\lambda$ strategy involves adding both children and parents to produce the next generation. This strategy was selected to maintain diversity in the population and prevent premature convergence to suboptimal solutions. In the context of this research, we set $\mu$ to 150 and $\lambda$ to 150. This configuration, chosen based on empirical testing, provides a balance between exploration of the search space and computational feasibility. Increasing $\mu$ and $\lambda$ beyond these values resulted in marginal improvements at significantly higher computational costs.
Initially, pairs of parents are chosen through tournament selection with a tournament size of 2. Subsequently, the children are generated employing a two-point crossover technique on the parents’ chromosomes with an 80% probability and applying a random resetting mutation with a 20% probability. The random resetting mutation implies that each gene of every child has a 20% chance of acquiring a new random value within its defined range. These probabilities were chosen based on preliminary experiments that demonstrated a good balance between exploration and exploitation. To create an offspring of 150 children, this selection, crossover, and mutation process is repeated 75 times. Finally, we combine the 150 children and the 150 parents, resulting in generations of 300 individuals, and select the top 150 ($\mu$) according to their fitness value to pass to the next generation.
The evolutionary process consists of 20 iterations, thereby generating a total of 3,000 individuals—each representing a distinct hyperparameter configuration. The choice of 20 iterations was determined based on convergence analysis, where we observed that fitness improvements plateaued after approximately 15–20 generations. From this pool of configurations, the one exhibiting the highest fitness is ultimately chosen as the optimal outcome.
@@ 5.2 Generating a default score with BERT
In this section, we delineate the methodology employed to generate a default score based on the textual description of the loan. We initiate the process by applying transfer learning utilizing an LLM, specifically BERT in our case. The fine-tuning of BERT, illustrated in Figure 2, results in a model that produces an outcome within the range of 0 to 1, offering a nuanced indicator rather than a binary classification. This subtle indicator is subsequently integrated into a classifier along with other input variables to predict the likelihood of loan default. The subsequent steps in this process are detailed next.
Figure 2: Fine-tuning process of BERT for generating a default score from loan descriptions.
@@ 5.2.1 Transfer learning to produce the default score
As outlined in Section 3, during the transfer learning phase, we aim to exploit BERT’s advanced language understanding while minimizing the need to learn from scratch. To achieve this, we freeze the weights of all but the last hidden layer of BERT. This approach preserves the model’s pre-trained capabilities and prevents overfitting on our dataset, while also mitigating catastrophic forgetting—a phenomenon where neural networks lose previously acquired knowledge when retrained on new tasks [7].
Subsequently, in the fine-tuning stage, only the last BERT layer and the newly added layers are adjusted to better serve our specific task of generating a default score. These layers are expected to enhance the model’s adaptability to our particular requirements, allowing slight parameter adjustments for improved task-specific performance, while maintaining the general language understanding gained from BERT’s initial pre-training. This strategy effectively balances specialized learning with the retention of valuable pre-trained knowledge.
To ensure that configurations are chosen based on empirical evidence, we incorporate architectural features into the training process. This allows the performance metric—in our case, balanced accuracy—to objectively guide model selection and minimize potential bias from manually imposed configurations. The parameters defining the extra layers’ architecture play a crucial role in balancing the model’s learning capacity and generalization ability—where, for instance, the second layer enhances learning, and the dropout layer promotes generalization. Thus, we explore various parameter configurations to determine the one that produces the best results. Specifically, we explore the 126 configurations that result from combining the following options:
@@
The neural network is trained to predict the loan outcome (default or non-default) based on the textual description as input. As a loss function, we use the weighted binary cross-entropy, which quantifies the difference between the predicted probabilities and the true binary labels by assigning different weights to each class. This approach ensures that the model does not bias predictions towards the majority class in imbalanced datasets, as it penalizes errors on the minority class more heavily. Considering class weights is crucial given the imbalanced nature of our dataset, where defaulted loans constitute only 15.27% of the total instances. The training was set up with a batch size of 64 and trained for 25 epochs, with an early stopper of 3 epochs.